### v1

# I’m Sorry Dave, I’m Afraid I Can’t Return That: On YouTube Search API Use in Research

Abstract
YouTube is among the most widely-used platforms worldwide, and has seen a lot of recent academic attention. Despite its popularity and the number of studies conducted on it, much less is understood about the way in which YouTube’s Data API, and especially the Search endpoint, operates. In this paper, we analyze the API’s behavior by running identical queries across a period of 12 weeks. Our findings suggest that the search endpoint returns highly inconsistent results between queries in ways that are not officially documented. Specifically, the API seems to randomize returned videos based on the relative popularity of the respective topic during the query period, making it nearly impossible to obtain representative historical video samples, especially during non-peak topical periods. Our results also suggest that the API may prioritize shorter, more popular videos, although the role of channel popularity is not as clear. We conclude with suggested strategies for researchers using the API for data collection, as well as future research directions on expanding the API’s use-cases.
1 Introduction
YouTube has been used to study a wide array of crucial societal problems like online hate [1, 17], accessibility [13], pseudoscientific misinformation [18], online scams [10], and child exposure to inappropriate content [5, 9, 16]. Much of this work has made use of the YouTube Data API, which offers several endpoints.
The way we understand these problems is influenced by the quality of the data we can get from the platform, yet researchers have pointed out that the YouTube API can behave inconsistently and return seemingly biased results [3, 24]. Beyond documentation and implementation errors [12], some works suggest that some of these biases may result from systematic API behavior, particularly through the keyword-based Search: list endpoint (“search” henceforth), that makes obtaining random samples very difficult [14, 22].
Although YouTube data are collected in a variety of ways, including crawling through recommended videos [11, 21, 23], deploying sockpuppet accounts [7, 8], extensions installed on participants’ browsers [6], or collecting videos from pre-curated channels [2, 9], the search endpoint remains an important part of the data collection pipeline for a lot of research [1, 13, 15, 16, 17, 19, 24]. Given its importance and the fact that its behavior is poorly understood [3, 14], we conduct an audit of this API endpoint. We run identical queries at 5-day intervals across a period of approximately 3 months to answer the following research questions:
  • RQ1
    How consistent and reliable are the data returned by the Search endpoint?
  • RQ2
    How does the API determine which videos are returned?
Through this work, we aim to inform better search strategies when using the YouTube API in terms of replicability and API token economy, while also exploring new ways in which the search endpoint can be used in academic research.
2 Background & Related Work
Based on the official documentation, [^1] https://developers.google.com/youtube/v3/docs/search/list the search endpoint allows a user to search by keywords, location, or live events, and enables filtering by several other parameters such as date ranges, specific channels, etc. However, this endpoint has a quota cost of 100 units per query; this is considerably higher than ID-based endpoints, which typically only cost 1 quota unit per query. With the default daily quota for a newly created client being 10,000, this only allows 100 search queries per day per client. However, the YouTube Data API has a researcher access program that provides substantially higher quotas to vetted accounts. This endpoint is not designed for volume, and the maximum number of results per query is 500 (a maximum of 50 per page and a maximum of 10 pages) [25].
One approach for comprehensive data collection was to identify topical “seed” channels and videos (either through external sources [16] or keyword search [18]), and obtain videos recommended by YouTube as being relevant to those initial sets to expand the dataset. However, YouTube deprecated the relatedToVideoId response that enabled this approach in 2023, effectively eliminating it from being conducted through the API.
An alternative advocated strategy has been time-split queries for clients endowed with sufficient quota [25]. The API allows the client to add parameters for publishedAfter and publishedBefore to restrict the data collection period. Researchers have used this to query in a “one per X time” fashion [22], where the observation period is split into time bins, each of which is queried separately to circumvent the 500-video limit [20, 24]. In theory, this should enable researchers to obtain every video uploaded on the topic, unless more than 500 videos were uploaded on a specific day.
Another commonly used strategy is identifying relevant channels through external sources like SocialBlade [9] or Reddit [2] and querying the API for their videos. While this can be done using several endpoints, for example, using Channels: list to extract a “playlist” of a channel’s uploaded videos and then querying the PlaylistItems: list endpoint for those videos, or adding a “channelId” parameter to the search endpoint, few papers clarify the exact endpoints used [13] or the dates on which the queries themselves are made [18]. As we later show, both of these can influence returned data.
3 Methods
For our experiments, we choose a diverse range of political, scientific, and entertainment topics that are either regional or international and vary in size and recency:
  • Black Lives Matter (BLM)
  • Brexit
  • US Capitol Riots
  • Grammy Awards (2024)
  • Higgs Boson
  • World Cup (2014)
For each topic, we run the same query every five days through the YouTube Data API v3 Search: list endpoint, starting on February 9 and ending on April 30, 2025 (see Appendix A for exact queries). Due to a technical problem, the collection on April 5th was skipped. In each query, we obtain a focal date (e.g., the day of the referendum for Brexit), and we set our data collection period between two weeks before and after this focal date (i.e., a total collection span of 28 days per topic). We then send queries for every hour within these 28 days to circumvent the maximum limits in returned videos imposed by the API following the strategy outlined in Section 2, resulting in 4032 total queries for every collection (24 hours $\times$ 28 days $\times$ 6 topics). This results in 16 snapshots over a period of 12 weeks. All queries are set to return results in reverse chronological order. We choose this order as it is an immutable video property, whereas other ordering options like view count or relevance may change over time. Thus, it offers the best baseline to study API consistency.
We then compare the consistency in the videos returned per collection using video ID set similarities. We find that this similarity decays over time, indicating that datasets collected using the exact same historical query may differ vastly based simply on when the queries were made. Moreover, we analyze the historical time points on which the most data is returned, finding that the API may be suppressing data returns for periods of relative topical inactivity even though returning these data would not bring the response above the maximum number of videos allowed by the API. Using a second-order Markov chain to model “transitions” between the presence or the absence of a video in successive collections, we find that video omission or inclusion is mostly conducted in a “rolling window” fashion. Finally, we analyze whether any video metadata (e.g., likes, views, etc.) is associated with more consistent video returns, finding that the API is more likely to return more popular videos that are drawn from less active topics.
4 API Behavior
In this section, we highlight how the API’s search endpoint returns inconsistent results. We document some of the potential mechanisms of this inconsistency and how videos may be drawn when determining what to return to the client.
4.1 Temporal Inconsistencies
For every collection instance at time $t$, we obtain the set of video IDs returned $S_{t}$ and calculate its Jaccard similarity with the set obtained in the previous collection $S_{t-1}$ and the very first collection $S_{t-n}$. We plot these rolling Jaccard similarities in Figure 1. In addition, we show descriptive statistics for the number of videos returned in Table 1.
Figure 1: Jaccard similarities of video ID sets relative to the previous and the first collection instance using the “search” endpoint. “Error bars” represent the set difference of $S_{t-1}-S_{t}$ (bottom bars) and $S_{t}-S_{t-1}$ (top bars).
topic min max mean std
BLM 639 765 743.44 27.86
Brexit 478 573 559.81 21.86
Capitol 507 590 571.81 17.35
Grammys 564 677 659.13 25.45
Higgs 476 512 507.44 8.32
World Cup 419 516 502.5 21.96
Table 1: Descriptive statistics for number of videos returned per topic across collections.
We find some non-constant differences between successive runs. However, over time, these differences compound to form video sets that are vastly different from the initial collection, to Jaccard values as low as $\sim$0.3 after 3 months. This equates to only 46% of the videos per set being shared, with the majority being unshared. The exception is the Higgs topic, which retains much higher consistency than the rest–we offer an explanation of why that may be in Section 5. As shown by the “error bars” in Figure 1, content deletions cannot explain this phenomenon as we find videos at $t$ that were not seen at $t-x$, despite these queries being fully historical (i.e., not spanning the query date). Moreover, we verify that this behavior does not extend to other endpoints that use ID-based queries (see Appendix B).
4.2 Randomization Mechanisms
An obvious question is whether the discrepancies are due to ceiling effects, despite the query specifying that results should be returned in chronological order. That is, if the eligible videos for a particular query exceed the maximum of 50 per page/500 per query allowed by the API, then we might expect that the results returned will be randomized. In that case, we should expect that some of our queries return more than 500 eligible videos, and that Jaccard similarities between collections are lower for hours on which more videos are uploaded. In this section, we test this theory.
We first obtain some descriptive statistics of videos returned for each hour and each of the topics (Table 2). The maximum number of videos returned for any given hour (max) remains well below the theoretical maximum of 50 per page, ruling out the ceiling effect explanation. Moreover, we compute the Spearman coefficient $\rho$ for the correlation between the Jaccard similarity of sets $T_{1}$ (first collection) and $T_{L}$ (last collection) and the average number of videos returned for that hour, after dropping all hours for which 0 videos are returned across collections to avoid inflating Jaccard similarity values based on empty sets. Based on the above reasoning, we should expect negative correlations between these variables. However, we observe almost the exact opposite pattern: For all but the Higgs topic, for which we observe a non-significant (negative) correlation, there are weak positive correlations between the number of videos for that hour and the Jaccard similarity, indicating that similarity values are, on average, higher for busier hours. Although this may be an artifact of the idea that more videos simply stabilize the Jaccard value, this analysis demonstrates that fewer videos do not necessarily mean that those videos will be the same across collections.
topic mean min max std $\rho$ N
BLM 1.10 0 17 2.33 **0.13 267
Brexit 0.83 0 13 1.57 ***0.15 324
Capitol 0.85 0 28 2.54 ***0.29 242
Grammys 0.98 0 21 2.22 ***0.26 387
Higgs 0.75 0 14 1.62 -0.11 216
World Cup 0.75 0 31 1.37 *0.12 418
Table 2: Descriptive statistics for per-hour number of videos returned. *p ¡ 0.05, **p ¡ 0.01, ***p ¡ 0.001. N is the number of videos retained after all hours with no videos returned across collections are dropped.
We also plot the daily video frequencies of the first and last collections alongside the Jaccard similarity for these daily sets in Figure 2. To ensure that these are not edge cases, we further plot the average daily frequencies across all collections. As can be seen, the average daily frequency distributions per collection map almost perfectly on each other. However, the volume of videos returned does not map onto the Jaccard similarities in any consistent manner, confirming our above findings of weak or non-existent correlations between these two factors. Most videos returned are uploaded around the focal date, with the exception of the BLM topic (likely due to when the protests surrounding George Floyd’s death intensified; the topical peak is recorded on Blackout Tuesday). Jaccard similarities at frequency peaks are comparable to other days on which much fewer videos are uploaded, indicating that the YouTube Data API operates on time-dependent systematic randomization. Interestingly, the YouTube recommendation system uses an empirical distribution of video popularity against its age to determine how relevant it is to recommend to users [4]. These patterns suggest that the API may similarly sample from topic-wide empirical distributions in terms of over-time interest.
Figure 2: Daily frequencies of videos returned with daily Jaccard similarities between first and last collections. Dashed vertical lines represent the corresponding topic’s D-day.
Overall, we do not find evidence of ceiling effects driving the API discrepancies. Instead, our results suggest that the YouTube API samples videos from empirical distributions, returning results based on the relative density of topical interest and even forcing zero videos to be returned when this relative density is adequately low. However, it is unclear how this interest is computed (e.g., if it is the volume of videos uploaded or something else).
4.3 Attrition Analysis
Next, we focus on whether this sampling operates on a “drop-in/drop-out” basis; that is, are videos more likely to reappear or remain left out in successive collections? To answer this, we utilize a second-order Markov chain where we treat the presence (P) or absence (A) of a video in any given collection as the two possible states. Then, across all topics and videos, we compute the transition probability from the two most recent states to the next one in a sliding window. We show the resulting transition probabilities in Figure 3.
Figure 3: Transition probabilities for presence or absence of videos in a second-order Markov chain. Columns show outgoing probabilities to the rows.
The results suggest that drop-ins and drop-outs are the normative behavior. That is, a video is more likely to be present or absent in a collection if it is present or absent, respectively, in the immediately previous collection. Moreover, this probability is higher when both previous states are the same. Therefore, the probability that a video is returned in the collection set may be influenced not only by the video’s upload date, but by the request date itself and whether the video is in the “windowed set” for that date.
5 Factors Behind Return Likelihood
Our next analysis focuses on whether the YouTube API is more likely to return videos with certain features at higher rates. For each video per topic, we count the number of times it is returned in our API calls, and we use this frequency as our dependent variable. Frequency ranges from 1-16 (contingent on our number of collections), and we split it into four roughly equal bins (1-5, 6-10, 11-15, 16), taking into account that 16 is the modal value . We then perform an ordinal regression using a logit link function (hence the binning). For robustness, we also perform an unbinned ordinal regression and a multiple regression taking frequency as a continuous variable; we report these models in Appendix C and find largely consistent results. As candidate predictors, we use the video-level features of topic, video duration, video definition, number of views, comments, and likes; and the channel-level features of channel views, subscribers, and number of videos uploaded. All continuous features are log-transformed to reduce multicollinearity and standardized for better comparison between coefficients.
A log-likelihood test shows that this model significantly outperforms the null model ($\chi^{2}=1137.63,p<0.001)$, although the overall fit is low (pseudo-$R^{2}=0.079$). This suggests either that several other factors may influence video appearance, or that much of the variance is indeed random. We show the standardized beta coefficients and confidence measures in Table 3. Topics are compared against BLM as the reference category, and the standard quality is compared against HD.
Variable $\beta$ SE 95% CI
SD (quality) -0.018 0.079 [-0.173, 0.137]
brexit (topic) ***1.231 0.098 [1.039, 1.423]
capriot (topic) -0.160 0.093 [-0.341, 0.022]
grammys (topic) *0.171 0.083 [0.008, 0.333]
higgs (topic) ***3.10 0.141 [2.826, 3.379]
worldcup (topic) 0.161 0.101 [-0.037, 0.359]
duration ***-0.115 0.028 [-0.170, -0.061]
views 0.161 0.088 [-0.011, 0.333]
likes **0.285 0.095 [0.098, 0.471]
comments 0.069 0.064 [-0.058, 0.195]
channel age 0.049 0.031 [-0.012, 0.110]
channel views *0.3176 0.135 [0.053, 0.582]
channel subs **-0.3784 0.122 [-0.617, -0.140]
# channel videos -0.0212 0.075 [-0.169, 0.126]
Table 3: Standardized regression coefficients for binned ordinal model. *$p<0.05$, **$p<0.01$, ***$p<0.001$.
Starting with video features, we see a negative effect of video duration and a positive effect of number of likes, indicating that shorter, more widely-liked videos are returned in more collections. We should note that the non-significant effect of views and comments is likely due to their very high correlations and thus shared variance with the number of likes ($r=0.92,p<0.001$ and $r=0.89,p<0.001$, respectively). Indeed, these predictors become significant when likes are dropped from the model, suggesting that more popular videos overall are returned more often. For channel features, the total number of channel views also has a positive effect, while the number of channel subscribers has a negative one. Curiously, despite a near-perfect correlation between these two variables ($r=0.97,p<0.001$), the negative effect of channel subscriptions persists when total channel views are dropped; however, total channel views become non-significant when subscriptions are dropped. For this reason, we urge caution in interpreting channel-related results as they may be spurious.
With respect to topics, we find that the Brexit, Grammy Awards, and Higgs Boson topics are returned in significantly more collections than BLM. Capitol Riot and World Cup show no effects. Although these topics look unrelated at first, a deeper look reveals that they are linked by one factor: size. We determine this using metadata returned with every query we send. As part of this metadata, the YouTube API returns a value called pageInfo.totalResults which reflects “the total number of results in the result set”, [^2] https://developers.google.com/youtube/v3/docs/search/list i.e., the total number of videos that match our query, with a maximum potential value of 1,000,000. We obtain the minimum, maximum, mean, and modal values from this field across every hour and collection run that we query per topic. As seen in Table 4, the three topics whose videos have higher appearance frequencies are also the smallest (and the only ones without a modal value of 1M, which is the maximum). Although this observation requires further experiments to verify, the results indicate that queries with a smaller video pool to draw from may return more consistent results (and may explain why Higgs, which is smaller by orders of magnitude compared to other topics, is also by far the most consistent in Figure 1).
Topic Min Max Mean Mode
BLM 679k 1M 982k 1M
Brexit 247k 786k 624k 613k
Capitol Riot 515k 1M 966k 1M
Grammys 12.8k 1M 150k 123k
Higgs 5.50k 65.2k 40.2k 39.0k
World Cup 634k 1M 998k 1M
Table 4: Potential video pool size per topic.
It is also interesting to note that, despite the modal number of videos returned across all collection hours for all topics being 0, the modal value for the available pool is much higher (and the maximum of 1M for 3/6 topics). Since it is unlikely that 1M videos are consistently uploaded every hour for any of these topics, this behavior suggests that the API does not take into account time constraints in determining the total pool of available videos.
Comparing Tables 1 and  4, it is also striking that the number of videos returned is much closer across topics than their respective topic sizes suggest, which is consistent with our distribution density explanation. This becomes even more apparent when scrutinizing the y-axis on Figure 2: The most-populated peaks are recorded for topics where the rest of the time-series is relatively inactive (e.g., Capitol Riot, Grammys), while topics that are more active throughout (e.g., World Cup, which is an ongoing tournament rather than a one-off event) record peaks at lower absolute values, most likely due to the number of videos to be returned being fixed and the actual videos being drawn from an empirical distribution.
6 Discussion
In this short paper, we contribute to our understanding of a crucial research tool. The behaviors we record offer both new strategies of working with the YouTube API, as well as opening further research directions to better understand how researchers can use it.
6.1 Implications
Our findings show that binning queries across the observation period is not as fruitful as previously thought [22, 24, 25], and offers low return on investment considering the quota cost of the search endpoint. Instead, researchers may experiment with breaking up their topics as opposed to their time frames. This can be achieved by incorporating more AND statements in the query or querying for multiple sub-topics (e.g., specific players alongside their national teams instead of the entirety of the World Cup event). The total number of results in the query metadata is a crucial way of assessing how optimal a query is (with lower being better/more stable).
Moreover, researchers who use the API for data collection should aim to use ID-based endpoints wherever possible, as these are more complete and reliable. The strategy of pre-selecting channels, provided that there are ways to obtain an adequately representative sample, is a viable one as long as the search endpoint is not used to collect their videos. Instead, as we outline in Section 2, researchers can use a combination of the Channels: list and PlaylistItems: list endpoints, both of which take IDs as queries, to get complete channel uploads.
Lastly, we urge researchers to clearly outline which endpoints they use in their data collection pipelines, as these choices can massively impact the replicability of their work.
6.2 Limitations and Future Work
The new strategies we put forward may need experimental validation, for example, by running progressively more restrictive queries and seeing how that influences the replicability of the data returned (alongside the reported video pool size).
Moreover, we do not explicitly analyze how other parameters, such as ordering or including channel IDs, may affect replicability, although we have no reason to believe that this would alter our conclusions (especially as reverse chronological order should be immutable for historical queries, other than video deletions).
Finally, there are several future directions that scholars can pursue to better understand this invaluable tool. For example, future research can replicate our experiments with more sparse collections over a longer period, to check for potential periodicity in set similarities. Moreover, given the substantial efforts that scholars have expended in creating sockpuppets for YouTube SERP audits [7, 8], future research can employ similar methods to ours to check the consistency between results of sockpuppet SERPs and search endpoint results. This would help us understand if the search endpoint has research value beyond data collection, for example, as a low-resource way of conducting SERP audits.
Acknowledgements. This work has been supported by the University of Washington’s Center for an Informed Public, the John S. and James L. Knight Foundation (G-2019-58788), and the William and Flora Hewlett Foundation (2023-02789). The author is grateful to Adrian Padilla, Bernhard Rieder, and Ethan Zuckerman for feedback, comments, and discussions related to this work.
Appendix A Query Parameters
Our general query parameters, followed by topic-specific parameters (keywords and dates).
A.1 General parameters
Unless [variable], these parameters were kept consistent across queries.
        "part": "snippet",
        "maxResults": 50,
        "order": "date",
        "safeSearch": "none",
        "publishedAfter": [variable],
        "publishedBefore": [variable],
        "type": "video",
        "q": [variable]
A.2 Topic-specific parameters
Keywords (q) and dates queried. Note that, for dates, we passed “publishedAfter” as the topic-specific date -14 and “publishedBefore” as +14 days.
Black Lives Matter
Focal date: Killing of George Floyd.
        "date": "2020-05-25T00:00:00Z",
        "q": "black lives matter"
Brexit
Focal date: Day of the referendum.
        "date": "2016-06-23T00:00:00Z"
        "q": "brexit referendum"
Capitol riots
Focal date: January 6th attack on the US Capitol.
        "date": "2021-01-06T00:00:00Z",
        "q": "us capitol"
Grammys 2024
Focal date: Day of the Awards ceremony.
        "date": "2024-02-04T00:00:00Z",
        "q": "grammy awards"
Higgs Boson
Focal date: Announcement of the “God particle” discovery.
        "date": "2012-07-04T00:00:00Z",
        "q": "higgs boson"
World Cup 2014
Focal date: Start and first game of the tournament.
        "date": "2014-06-12T00:00:00Z",
        "q": "fifa world cup"
Appendix B ID-Based Queries
This section covers tests conducted with API endpoints that accept video or other IDs as queries. These endpoints show stable behavior and return mostly consistent data.
B.1 Video: list Endpoint
In Figure 4, we show common video IDs between collections at a given time t and the previous time, as well as the first collection. For each collection, we query the Video: list endpoint immediately after obtaining results through the search endpoint to get details and metadata about videos, such as their descriptions, view counts, likes, etc., using the video IDs. We compute the percentage of videos for which metadata is returned at t and t-1, and we also obtain Jaccard similarities for the videos returned between $S_{t}$ and $S_{t-1}$, as well as between $S_{t}$ and $S_{1}$. Since these comparisons are restricted only to video IDs that are common in both sets being compared, the overall coverage and Jaccard similarity are higher for this endpoint. Moreover, the fact that we do not find consistent patterns between comparison ID and $J(S_{t},S_{1})$ suggests that API gaps in returning specific video metadata are not systematic, and are thus likely errors rather than intentional API behavior.
Figure 4: Parallel plots of percentage of common videos retrieved between successive runs and overall Jaccard similarity of common videos retrieved using the “Videos: list” endpoint. Comparison ID refers to the respective collection (higher ID = later collection).
B.2 CommentThreads: list and Comments: list Endpoints
We also query the CommentThreads: list endpoint, which accepts video IDs as queries and returns all comment threads (with a maximum of five nested comments), as well as the Comments: list endpoint, which accepts thread IDs as queries and returns all nested comments. Due to the scale and number of comments and comment threads returned in each instance, we only make comparisons between the first and last collection. Moreover, we only consider comments that were posted at most 3 weeks after the given topic’s D-day (we allow an additional week beyond our video collection stopping point to allow for consolidation of comments on videos that were uploaded later). Given the findings presented in Figure 1, we assume that any differences will be maximized the farther apart the collections occur.
For each collection instance, we compare the similarities of the set of top-level and nested comments returned across both all videos returned in that respective collection, and across videos that are common in both collections. We show these results in Table 5. Unsurprisingly, we find some deviations between both top-level and nested comment sets between the first and last collections, as these are drawn from different parent videos. However, the deviation patterns are not necessarily consistent with the magnitude of deviations in video IDs themselves. For example, although the Brexit topic shows overall lower video deviations than other topics (except for Higgs), it shows the second-highest deviation behind only BLM in both top-level and nested comments; this may be an artifact of higher activity under more contested topics. With respect to comments drawn from common videos, differences between both nested and top-level comments between collections are negligible, showing that this endpoint itself does not systematically randomize results and is likely returning (almost) all comments for every queried video.
topic TL, NS N, NS TL, S N, S
BLM .329 .307 .976 .983
Brexit .381 .339 .999 .999
Capitol .648 .625 .998 .994
Grammys .728 .737 .996 .992
Higgs .974 N/A .998 N/A
World Cup .470 .532 .999 .999
Table 5: Jaccard similarities between first- and last-collection comment sets. TL = top-level, N = nested. NS = non-shared videos (full sets), S = shared videos only. N/A values for the Higgs topic, which is the oldest, are possibly due to the comment reply affordance operating differently in 2012.
Appendix C Regression Robustness Checks
Alternative regression model setups as robustness checks against the binned regression model implemented in the main paper.
C.1 Frequency as Continuous Variable
We use frequency as our dependent variable in a multiple Ordinary Least Squares (OLS) regression with robust standard errors. The overall model is significant ($F_{(14,5348)}=122.3,p<0.001$) and shows modest fit ($R^{2}=0.164$). We report standardized beta coefficients with confidence metrics in Table 6. The patterns are identical to what is reported in the main paper.
Variable $\beta$ SE 95% CI
SD (quality) 0.0712 0.205 [-0.331, 0.474]
brexit (topic) ***3.416 0.274 [2.878, 3.953]
capriot (topic) -0.283 0.257 [-0.786, 0.220]
grammys (topic) *0.571 0.238 [0.105, 1.038]
higgs (topic) ***6.718 0.248 [6.231, 7.205]
worldcup (topic) 0.438 0.288 [-0.126, 1.003]
duration ***-0.285 0.076 [-0.435, -0.135]
views 0.429 0.238 [-0.037, 0.896]
likes **0.713 0.262 [0.198, 1.227]
comments 0.242 0.177 [-0.105, 0.588]
channel age 0.113 0.084 [-0.052, 0.279]
channel views **1.079 0.349 [0.394, 1.763]
channel subs ***-1.157 0.319 [-1.783, -0.531]
# channel videos -0.2212 0.208 [-0.629, 0.187]
Table 6: Standardized regression coefficients for OLS model. *$p<0.05$, **$p<0.01$, ***$p<0.001$.
C.2 Non-Binned Ordinal Regression
Ordinal regression where frequencies are treated as 16 distinct categories. We use a complementary log-log link function instead of logit due to the distribution being skewed towards the highest value. The overall model performs significantly better against a null model ($\chi^{2}=1167.64,P<0.001$), although the overall fit is low (pseudo-$R^{2}=0.04$). Coefficients are reported in Table 7. Patterns are largely consistent with the other models, except for the World Cup topic now also showing marginally significant differences (higher return frequencies) compared to BLM.
Variable $\beta$ SE 95% CI
SD (quality) 0.0228 0.051 [-0.077, 0.122]
brexit (topic) ***0.9207 0.065 [0.793, 1.049]
capriot (topic) -0.0412 0.059 [-0.156, 0.074]
grammys (topic) ***0.2395 0.051 [0.139, 0.340]
higgs (topic) ***2.2998 0.115 [2.075, 2.525]
worldcup (topic) *0.1338 0.066 [0.004, 0.264]
duration ***-0.0710 0.018 [-0.106, -0.036]
views 0.0352 0.056 [-0.074, 0.145]
likes **0.2051 0.062 [0.084, 0.326]
comments 0.0656 0.042 [-0.017, 0.148]
channel age 0.0355 0.019 [-0.002, 0.073]
channel views **0.2852 0.093 [0.103, 0.468]
channel subs **-0.2734 0.081 [-0.431, -0.116]
# channel videos -0.0958 0.049 [-0.193, 0.001]
Table 7: Standardized regression coefficients for full ordered model. *$p<0.05$, **$p<0.01$, ***$p<0.001$.

### v2

@@
-I’m Sorry Dave, I’m Afraid I Can’t Return That: On YouTube Search API Use in Research
-Abstract
+On YouTube Search API Use in Research
+Abstract.
 YouTube is among the most widely-used platforms worldwide, and has seen a lot of recent academic attention.
@@
 In this paper, we analyze the API’s behavior by running identical queries across a period of 12 weeks.
-Our findings suggest that the search endpoint returns highly inconsistent results between queries in ways that are not officially documented.
+Our findings show that the search endpoint returns highly variable results between queries.
 Specifically, the API seems to randomize returned videos based on the relative popularity of the respective topic during the query period, making it nearly impossible to obtain representative historical video samples, especially during non-peak topical periods.
@@
 We conclude with suggested strategies for researchers using the API for data collection, as well as future research directions on expanding the API’s use-cases.
-1 Introduction
-YouTube has been used to study a wide array of crucial societal problems like online hate [1, 17], accessibility [13], pseudoscientific misinformation [18], online scams [10], and child exposure to inappropriate content [5, 9, 16].
+YouTube, API, Research tools, Data collection
+journalyear: 2025copyright: ccconference: Proceedings of the 2025 ACM Internet Measurement Conference; October 28–31, 2025; Madison, WI, USAbooktitle: Proceedings of the 2025 ACM Internet Measurement Conference (IMC ’25), October 28–31, 2025, Madison, WI, USAdoi: 10.1145/3730567.3764492isbn: 979-8-4007-1860-1/2025/10ccs: Information systems Information retrievalccs: Information systems World Wide Webccs: Human-centered computing HCI design and evaluation methodsccs: Human-centered computing Collaborative and social computing
+1. Introduction
+YouTube has been used to study a wide array of crucial societal problems like online hate (Albadi et al., 2022; Papadamou et al., 2021), accessibility (May et al., 2024), pseudoscientific misinformation (Papadamou et al., 2022), online scams (Liu et al., 2024), and child exposure to inappropriate content (Gkolemi et al., 2022; Khan et al., 2024; Papadamou et al., 2020).
 Much of this work has made use of the YouTube Data API, which offers several endpoints.
-The way we understand these problems is influenced by the quality of the data we can get from the platform, yet researchers have pointed out that the YouTube API can behave inconsistently and return seemingly biased results [3, 24].
-Beyond documentation and implementation errors [12], some works suggest that some of these biases may result from systematic API behavior, particularly through the keyword-based Search: list endpoint (“search” henceforth), that makes obtaining random samples very difficult [14, 22].
-Although YouTube data are collected in a variety of ways, including crawling through recommended videos [11, 21, 23], deploying sockpuppet accounts [7, 8], extensions installed on participants’ browsers [6], or collecting videos from pre-curated channels [2, 9], the search endpoint remains an important part of the data collection pipeline for a lot of research [1, 13, 15, 16, 17, 19, 24].
-Given its importance and the fact that its behavior is poorly understood [3, 14], we conduct an audit of this API endpoint.
+However, researchers have pointed out that the YouTube API may sometimes return seemingly biased results (Choi et al., 2025; Violot et al., 2024).
+Beyond documentation and implementation errors (Martin-Lopez et al., 2021), some works suggest that this can result from systematic API behavior, particularly through the keyword-based Search: list endpoint (“search” henceforth), which makes obtaining random samples difficult (McGrady et al., 2023; Rieder et al., 2018).
+Although YouTube data are collected in a variety of ways, including crawling through recommended videos (Mamié et al., 2021; Ribeiro et al., 2020; Röchert et al., 2020), deploying sockpuppet accounts (Hussein et al., 2020; Jung et al., 2025), extensions installed on participants’ browsers (Hu et al., 2024), or collecting videos from pre-curated channels (Bertaglia et al., 2024; Khan et al., 2024), the search endpoint remains an important part of the data collection pipeline for a lot of research (Albadi et al., 2022; May et al., 2024; Nigatu and Raji, 2024; Papadamou et al., 2020, 2021; Pera and Aiello, 2024; Violot et al., 2024).
+Given its importance and the fact that its behavior is poorly understood (Choi et al., 2025; McGrady et al., 2023), we conduct an audit of this API endpoint.
 We run identical queries at 5-day intervals across a period of approximately 3 months to answer the following research questions:
   • RQ1
-    How consistent and reliable are the data returned by the Search endpoint?
+    How consistent are the data returned by the Search endpoint?
   • RQ2
     How does the API determine which videos are returned?
+Comparing the set similarities of videos obtained with each query, we find that this similarity decays over time, indicating that datasets collected using the exact same historical query may differ vastly based simply on when the queries were made.
+Moreover, we analyze the historical time points on which the most data is returned.
+We find that the API may be withholding data for periods of relative topical inactivity even though returning these data would not bring the response above the maximum number of videos allowed by the API.
+Using a second-order Markov chain to model “transitions” between the presence or the absence of a video in successive collections, we find that video omission or inclusion is mostly conducted in a “rolling window” fashion.
+Finally, we analyze whether any video metadata (e.g., likes, views, etc.) are associated with more consistent video returns, finding that the API is more likely to return more popular videos that are drawn from less active topics.
 Through this work, we aim to inform better search strategies when using the YouTube API in terms of replicability and API token economy, while also exploring new ways in which the search endpoint can be used in academic research.
-2 Background & Related Work
+2. Background
 Based on the official documentation, [^1] https://developers.google.com/youtube/v3/docs/search/list the search endpoint allows a user to search by keywords, location, or live events, and enables filtering by several other parameters such as date ranges, specific channels, etc.
-However, this endpoint has a quota cost of 100 units per query; this is considerably higher than ID-based endpoints, which typically only cost 1 quota unit per query.
-With the default daily quota for a newly created client being 10,000, this only allows 100 search queries per day per client.
-However, the YouTube Data API has a researcher access program that provides substantially higher quotas to vetted accounts.
-This endpoint is not designed for volume, and the maximum number of results per query is 500 (a maximum of 50 per page and a maximum of 10 pages) [25].
-One approach for comprehensive data collection was to identify topical “seed” channels and videos (either through external sources [16] or keyword search [18]), and obtain videos recommended by YouTube as being relevant to those initial sets to expand the dataset.
-However, YouTube deprecated the relatedToVideoId response that enabled this approach in 2023, effectively eliminating it from being conducted through the API.
-An alternative advocated strategy has been time-split queries for clients endowed with sufficient quota [25].
+However, this endpoint has a quota cost of 100 units per query; this is considerably higher than ID-based endpoints, which typically only cost 1 unit.
+With the default daily quota being 10,000, this allows 100 search queries per day per client.
+However, the YouTube Data API has a researcher access program that allows higher quotas to vetted accounts.
+This endpoint is not designed for volume.
+The maximum number of results per query is 500 (max.
+50 per page and max.
+10 pages) (Yin and Brown, 2018), thus, it is necessary for YouTube to sample the videos it returns.
+One approach to collect all videos on a topic was to identify topical “seed” channels and videos (either through external sources (Papadamou et al., 2020) or keyword search (Papadamou et al., 2022)), and obtain videos recommended by YouTube as being relevant to those initial sets to expand the dataset.
+However, the relatedToVideoId response that enabled this approach was deprecated in 2023, eliminating it from being conducted through the API.
+An alternative advocated strategy has been time-split queries for clients endowed with sufficient quota (Yin and Brown, 2018).
 The API allows the client to add parameters for publishedAfter and publishedBefore to restrict the data collection period.
-Researchers have used this to query in a “one per X time” fashion [22], where the observation period is split into time bins, each of which is queried separately to circumvent the 500-video limit [20, 24].
+Researchers have used this to query in a “one per X time” fashion (Rieder et al., 2018), where the observation period is split into time bins, each of which is queried separately to circumvent the 500-video limit (Porreca et al., 2020; Violot et al., 2024).
 In theory, this should enable researchers to obtain every video uploaded on the topic, unless more than 500 videos were uploaded on a specific day.
-Another commonly used strategy is identifying relevant channels through external sources like SocialBlade [9] or Reddit [2] and querying the API for their videos.
-While this can be done using several endpoints, for example, using Channels: list to extract a “playlist” of a channel’s uploaded videos and then querying the PlaylistItems: list endpoint for those videos, or adding a “channelId” parameter to the search endpoint, few papers clarify the exact endpoints used [13] or the dates on which the queries themselves are made [18].
+However, recent work finds a strong recency bias with this approach, with a much higher volume of videos for dates closer to the query date than for historical dates (Rieder et al., 2025).
+Another commonly used strategy is identifying relevant channels through external sources like SocialBlade (Khan et al., 2024) or Reddit (Bertaglia et al., 2024) and querying the API for their videos.
+While this can be done using several endpoints, for example, using Channels: list to extract a “playlist” of a channel’s uploaded videos and then querying the PlaylistItems: list endpoint for those videos, or adding a “channelId” parameter to the search endpoint, few papers clarify the exact endpoints used (May et al., 2024) or the dates on which the queries themselves are made (Papadamou et al., 2022).
 As we later show, both of these can influence returned data.
-3 Methods
-For our experiments, we choose a diverse range of political, scientific, and entertainment topics that are either regional or international and vary in size and recency:
-  • Black Lives Matter (BLM)
-  • Brexit
-  • US Capitol Riots
+3. Methods
+For our experiments, we choose a diverse range of political, scientific, and entertainment topics that are either regional or international and vary in size and recency.
+Although these topics are not exhaustive, they allow us to observe whether certain patterns may be due to characteristics like topic sensitivity.
+The exact queries per topic are shown in Appendix B:
+  • Black Lives Matter (BLM) (2020)
+  • Brexit (2016)
+  • US Capitol Riots (2021)
   • Grammy Awards (2024)
-  • Higgs Boson
+  • Higgs Boson (2012)
   • World Cup (2014)
-For each topic, we run the same query every five days through the YouTube Data API v3 Search: list endpoint, starting on February 9 and ending on April 30, 2025 (see Appendix A for exact queries).
+Each topic has a focal “D-day” on which a central event took place (e.g., the day of the referendum for Brexit; see Appendix B).
+We set our data collection period between two weeks before and after this date (i.e., a total collection span of 28 days per topic).
+For each topic, we run the same query every five days through the YouTube Data API v3 Search: list endpoint, starting on February 9 and ending on April 30, 2025.
 Due to a technical problem, the collection on April 5th was skipped.
-In each query, we obtain a focal date (e.g., the day of the referendum for Brexit), and we set our data collection period between two weeks before and after this focal date (i.e., a total collection span of 28 days per topic).
-We then send queries for every hour within these 28 days to circumvent the maximum limits in returned videos imposed by the API following the strategy outlined in Section 2, resulting in 4032 total queries for every collection (24 hours $\times$ 28 days $\times$ 6 topics).
-This results in 16 snapshots over a period of 12 weeks.
-All queries are set to return results in reverse chronological order.
+Our data and code are available on GitHub.
+[^2] https://github.com/alefstrat/youtube_api_audit
+We send queries for every hour within these 28 days to circumvent the maximum limits in returned videos imposed by the API following the strategy outlined in Section 2, resulting in 4032 total queries for every collection (24 hours $\times$ 28 days $\times$ 6 topics) and 16 snapshots over 12 weeks.
+Queries are made with an API token, not OAuth 2.0, so that account effects do not confound our results.
+Responses are set to be returned in reverse chronological order.
 We choose this order as it is an immutable video property, whereas other ordering options like view count or relevance may change over time.
 Thus, it offers the best baseline to study API consistency.
-We then compare the consistency in the videos returned per collection using video ID set similarities.
-We find that this similarity decays over time, indicating that datasets collected using the exact same historical query may differ vastly based simply on when the queries were made.
-Moreover, we analyze the historical time points on which the most data is returned, finding that the API may be suppressing data returns for periods of relative topical inactivity even though returning these data would not bring the response above the maximum number of videos allowed by the API.
-Using a second-order Markov chain to model “transitions” between the presence or the absence of a video in successive collections, we find that video omission or inclusion is mostly conducted in a “rolling window” fashion.
-Finally, we analyze whether any video metadata (e.g., likes, views, etc.) is associated with more consistent video returns, finding that the API is more likely to return more popular videos that are drawn from less active topics.
-4 API Behavior
-In this section, we highlight how the API’s search endpoint returns inconsistent results.
-We document some of the potential mechanisms of this inconsistency and how videos may be drawn when determining what to return to the client.
-4.1 Temporal Inconsistencies
+However, we stress that the YouTube API’s documentation makes no commitment to order results beyond a daily granularity.
+Our choice to make queries for each individual hour is so that we maintain more consistency between low-activity and high-activity days, i.e., avoiding ceiling effects for days on which more than 500 videos may have been posted.
+However, on such days, it can alternatively be the case that order may take precedence over time-filtering.
+That is, returned videos may be sampled from a given day, rather than a given hour, if the number of eligible results exceeds the maximum.
+4. API Behavior
+In this section, we highlight how the API’s search endpoint varies returned results.
+We document some of the potential mechanisms of this variability and how videos may be drawn when determining what to return to the client.
+4.1. Temporal Variability
 For every collection instance at time $t$, we obtain the set of video IDs returned $S_{t}$ and calculate its Jaccard similarity with the set obtained in the previous collection $S_{t-1}$ and the very first collection $S_{t-n}$.
 We plot these rolling Jaccard similarities in Figure 1.
-In addition, we show descriptive statistics for the number of videos returned in Table 1.
-Figure 1: Jaccard similarities of video ID sets relative to the previous and the first collection instance using the “search” endpoint.
+Table 1 shows descriptives for the total number of videos returned per collection.
+Figure 1.
+Jaccard similarities of video ID sets relative to the previous and the first collection instance using the “search” endpoint.
 “Error bars” represent the set difference of $S_{t-1}-S_{t}$ (bottom bars) and $S_{t}-S_{t-1}$ (top bars).
@@
 World Cup 419 516 502.5 21.96
-Table 1: Descriptive statistics for number of videos returned per topic across collections.
-We find some non-constant differences between successive runs.
-However, over time, these differences compound to form video sets that are vastly different from the initial collection, to Jaccard values as low as $\sim$0.3 after 3 months.
-This equates to only 46% of the videos per set being shared, with the majority being unshared.
+Table 1. Descriptive statistics for number of videos returned per topic across collections.
+We find differences between successive runs, which, over time, compound to form video sets that are vastly different from the initial collection.
 The exception is the Higgs topic, which retains much higher consistency than the rest–we offer an explanation of why that may be in Section 5.
-As shown by the “error bars” in Figure 1, content deletions cannot explain this phenomenon as we find videos at $t$ that were not seen at $t-x$, despite these queries being fully historical (i.e., not spanning the query date).
-Moreover, we verify that this behavior does not extend to other endpoints that use ID-based queries (see Appendix B).
-4.2 Randomization Mechanisms
-An obvious question is whether the discrepancies are due to ceiling effects, despite the query specifying that results should be returned in chronological order.
-That is, if the eligible videos for a particular query exceed the maximum of 50 per page/500 per query allowed by the API, then we might expect that the results returned will be randomized.
-In that case, we should expect that some of our queries return more than 500 eligible videos, and that Jaccard similarities between collections are lower for hours on which more videos are uploaded.
+The “error bars” in Figure 1 rule out content deletions as a potential explanation; we find videos at $t$ that were not seen at txt-x, despite these queries being fully historical (i.e., not spanning the query date).
+This behavior does not extend to endpoints that take IDs as queries (see Appendix C).
+4.2. Randomization Mechanisms
+An obvious question is whether the variability is due to ceiling effects, i.e., whether randomization occurs due to the possible matches exceeding the maximum number of results allowed to be returned.
 In this section, we test this theory.
@@
 The maximum number of videos returned for any given hour (max) remains well below the theoretical maximum of 50 per page, ruling out the ceiling effect explanation.
-Moreover, we compute the Spearman coefficient $\rho$ for the correlation between the Jaccard similarity of sets $T_{1}$ (first collection) and $T_{L}$ (last collection) and the average number of videos returned for that hour, after dropping all hours for which 0 videos are returned across collections to avoid inflating Jaccard similarity values based on empty sets.
-Based on the above reasoning, we should expect negative correlations between these variables.
+Moreover, we compute the Spearman coefficient $\rho$ for the correlation between the Jaccard similarity of sets $T_{1}$ (first collection) and $T_{L}$ (last collection) and the average number of videos returned for that hour.
+However, to avoid inflating Jaccard similarity values based on empty sets, we first drop all hours for which 0 videos are returned across collections.
+The correlation is meant as a soft test of the ceiling effect, since, if randomization is indeed more prominent when more videos are returned, we should expect lower Jaccard similarities with a higher number of returned videos (i.e., a negative correlation).
 However, we observe almost the exact opposite pattern: For all but the Higgs topic, for which we observe a non-significant (negative) correlation, there are weak positive correlations between the number of videos for that hour and the Jaccard similarity, indicating that similarity values are, on average, higher for busier hours.
@@
 World Cup 0.75 0 31 1.37 *0.12 418
-Table 2: Descriptive statistics for per-hour number of videos returned.
+Table 2.
+Descriptive statistics for per-hour number of videos returned.
 *p ¡ 0.05, **p ¡ 0.01, ***p ¡ 0.001.
 N is the number of videos retained after all hours with no videos returned across collections are dropped.
+Figure 2. Daily frequencies of videos returned with daily Jaccard similarities between first and last collections. Dashed vertical lines represent the corresponding topic’s D-day.
 We also plot the daily video frequencies of the first and last collections alongside the Jaccard similarity for these daily sets in Figure 2.
@@
 Jaccard similarities at frequency peaks are comparable to other days on which much fewer videos are uploaded, indicating that the YouTube Data API operates on time-dependent systematic randomization.
-Interestingly, the YouTube recommendation system uses an empirical distribution of video popularity against its age to determine how relevant it is to recommend to users [4].
+Interestingly, the YouTube recommendation system uses an empirical distribution of video popularity against its age to determine how relevant it is to recommend to users (Covington et al., 2016).
 These patterns suggest that the API may similarly sample from topic-wide empirical distributions in terms of over-time interest.
-Figure 2: Daily frequencies of videos returned with daily Jaccard similarities between first and last collections. Dashed vertical lines represent the corresponding topic’s D-day.
-Overall, we do not find evidence of ceiling effects driving the API discrepancies.
+Overall, we do not find evidence of ceiling effects driving the API randomization.
 Instead, our results suggest that the YouTube API samples videos from empirical distributions, returning results based on the relative density of topical interest and even forcing zero videos to be returned when this relative density is adequately low.
 However, it is unclear how this interest is computed (e.g., if it is the volume of videos uploaded or something else).
-4.3 Attrition Analysis
+4.3. Attrition Analysis
 Next, we focus on whether this sampling operates on a “drop-in/drop-out” basis; that is, are videos more likely to reappear or remain left out in successive collections?
@@
 We show the resulting transition probabilities in Figure 3.
-Figure 3: Transition probabilities for presence or absence of videos in a second-order Markov chain. Columns show outgoing probabilities to the rows.
+Figure 3.
+Transition probabilities for presence (P) or absence (A) of videos in a second-order Markov chain.
+Example interpretation: Column $A\rightarrow A$ and row $P$ show the probability of P given that the two previous states are A.
 The results suggest that drop-ins and drop-outs are the normative behavior.
@@
 Therefore, the probability that a video is returned in the collection set may be influenced not only by the video’s upload date, but by the request date itself and whether the video is in the “windowed set” for that date.
-5 Factors Behind Return Likelihood
+5. Factors Behind Return Likelihood
 Our next analysis focuses on whether the YouTube API is more likely to return videos with certain features at higher rates.
-For each video per topic, we count the number of times it is returned in our API calls, and we use this frequency as our dependent variable.
-Frequency ranges from 1-16 (contingent on our number of collections), and we split it into four roughly equal bins (1-5, 6-10, 11-15, 16), taking into account that 16 is the modal value .
-We then perform an ordinal regression using a logit link function (hence the binning).
-For robustness, we also perform an unbinned ordinal regression and a multiple regression taking frequency as a continuous variable; we report these models in Appendix C and find largely consistent results.
-As candidate predictors, we use the video-level features of topic, video duration, video definition, number of views, comments, and likes; and the channel-level features of channel views, subscribers, and number of videos uploaded.
-All continuous features are log-transformed to reduce multicollinearity and standardized for better comparison between coefficients.
-A log-likelihood test shows that this model significantly outperforms the null model ($\chi^{2}=1137.63,p<0.001)$, although the overall fit is low (pseudo-$R^{2}=0.079$).
-This suggests either that several other factors may influence video appearance, or that much of the variance is indeed random.
-We show the standardized beta coefficients and confidence measures in Table 3.
-Topics are compared against BLM as the reference category, and the standard quality is compared against HD.
-Variable $\beta$ SE 95% CI
-SD (quality) -0.018 0.079 [-0.173, 0.137]
-brexit (topic) ***1.231 0.098 [1.039, 1.423]
-capriot (topic) -0.160 0.093 [-0.341, 0.022]
-grammys (topic) *0.171 0.083 [0.008, 0.333]
-higgs (topic) ***3.10 0.141 [2.826, 3.379]
-worldcup (topic) 0.161 0.101 [-0.037, 0.359]
-duration ***-0.115 0.028 [-0.170, -0.061]
-views 0.161 0.088 [-0.011, 0.333]
-likes **0.285 0.095 [0.098, 0.471]
-comments 0.069 0.064 [-0.058, 0.195]
-channel age 0.049 0.031 [-0.012, 0.110]
-channel views *0.3176 0.135 [0.053, 0.582]
-channel subs **-0.3784 0.122 [-0.617, -0.140]
-# channel videos -0.0212 0.075 [-0.169, 0.126]
-Table 3: Standardized regression coefficients for binned ordinal model. *$p<0.05$, **$p<0.01$, ***$p<0.001$.
-Starting with video features, we see a negative effect of video duration and a positive effect of number of likes, indicating that shorter, more widely-liked videos are returned in more collections.
-We should note that the non-significant effect of views and comments is likely due to their very high correlations and thus shared variance with the number of likes ($r=0.92,p<0.001$ and $r=0.89,p<0.001$, respectively).
-Indeed, these predictors become significant when likes are dropped from the model, suggesting that more popular videos overall are returned more often.
-For channel features, the total number of channel views also has a positive effect, while the number of channel subscribers has a negative one.
-Curiously, despite a near-perfect correlation between these two variables ($r=0.97,p<0.001$), the negative effect of channel subscriptions persists when total channel views are dropped; however, total channel views become non-significant when subscriptions are dropped.
-For this reason, we urge caution in interpreting channel-related results as they may be spurious.
-With respect to topics, we find that the Brexit, Grammy Awards, and Higgs Boson topics are returned in significantly more collections than BLM.
-Capitol Riot and World Cup show no effects.
-Although these topics look unrelated at first, a deeper look reveals that they are linked by one factor: size.
-We determine this using metadata returned with every query we send.
-As part of this metadata, the YouTube API returns a value called pageInfo.totalResults which reflects “the total number of results in the result set”, [^2] https://developers.google.com/youtube/v3/docs/search/list i.e., the total number of videos that match our query, with a maximum potential value of 1,000,000.
+We treat this as an explainability problem where we count the number of times each video is returned in our API calls.
+We then determine the features that are most predictive of this frequency.
+Our candidate predictors are topic, video duration, video definition, number of views, comments, and likes at the video level; channel views, subscribers, and number of videos uploaded at the channel level.
+All continuous features are log-transformed to normalize their distributions.
+We deploy a gradient boosting model using LightGBM (Ke et al., 2017) with a Poisson regression objective, 100 estimators, and a learning rate of 0.1.
+The data are split into 80-20 train-test sets ($R^{2}=0.19)$.
+We examine the role of each feature from the predictions on the test set using SHapley Additive exPlanations (SHAP) (Lundberg and Lee, 2017), shown in Figure 4.
+Features are shown in descending order of importance.
+We also perform robustness checks using regression models with the same variables in Appendix D, which show directionally consistent results.
+Figure 4. Beeswarm plot of SHAP values per feature.
+Starting with video metadata, we see that the number of views tends to positively predict return frequency and is the most important feature.
+In terms of duration, which is also among the most important features, shorter videos tend to have more returns.
+There is also a slightly positive effect of likes, while the number of comments is not as important and video quality plays no role.
+Overall, this suggests that the API tends to return shorter, more popular videos.
+For channel features, returns are more likely from older channels with fewer videos.
+There is a slightly negative trend in terms of the number of channel subscriptions, while the role of accumulated number of channel views is more mixed.
+Overall, channel popularity is not as important, though other channel features like age and activity may play a role.
+Interestingly, the Higgs Boson and Brexit topics are highly important and positive features for return frequency; the Grammy’s topic also trends towards positive effects.
+The Capitol riots, BLM, and World Cup topics trend toward negative effects, though the latter two are not as important.
+Although the directionally grouped topics look unrelated at first, a deeper look reveals that they are linked by one factor: size.
+We determine this using metadata returned with every query we send, which contain a value called pageInfo.totalResults.
+This reflects the total number of results in the result set, i.e., the total number of videos that match our query, with a maximum potential value of 1,000,000.
 We obtain the minimum, maximum, mean, and modal values from this field across every hour and collection run that we query per topic.
 As seen in Table 3, the three topics whose videos have higher appearance frequencies are also the smallest (and the only ones without a modal value of 1M, which is the maximum).
-Although this observation requires further experiments to verify, the results indicate that queries with a smaller video pool to draw from may return more consistent results (and may explain why Higgs, which is smaller by orders of magnitude compared to other topics, is also by far the most consistent in Figure 1).
+The results indicate that queries with a smaller video pool to draw from may return more consistent results (and may explain why Higgs, which is by far the smallest topic, is also the most consistent in Figure 1), though this remains to be experimentally verified.
 Topic Min Max Mean Mode
@@
 World Cup 634k 1M 998k 1M
-Table 4: Potential video pool size per topic.
-It is also interesting to note that, despite the modal number of videos returned across all collection hours for all topics being 0, the modal value for the available pool is much higher (and the maximum of 1M for 3/6 topics).
-Since it is unlikely that 1M videos are consistently uploaded every hour for any of these topics, this behavior suggests that the API does not take into account time constraints in determining the total pool of available videos.
-Comparing Tables 1 and  4, it is also striking that the number of videos returned is much closer across topics than their respective topic sizes suggest, which is consistent with our distribution density explanation.
+Table 3. Potential video pool size per topic.
+Notably, despite the modal number of videos returned across all collection hours for all topics being 0, the modal value for the available pool is much higher (and the maximum of 1M for 3/6 topics).
+Since it is unlikely that 1M videos are consistently uploaded every hour for any of these topics, this suggests that time constraints in queries do not affect the total pool of available videos in the API’s results set.
+Rather, they may only filter the response after results are returned.
+Comparing Tables 1 and  3, it is also striking that the number of videos returned is much closer across topics than their respective topic sizes suggest, which is consistent with our distribution density explanation.
 This becomes even more apparent when scrutinizing the y-axis on Figure 2: The most-populated peaks are recorded for topics where the rest of the time-series is relatively inactive (e.g., Capitol Riot, Grammys), while topics that are more active throughout (e.g., World Cup, which is an ongoing tournament rather than a one-off event) record peaks at lower absolute values, most likely due to the number of videos to be returned being fixed and the actual videos being drawn from an empirical distribution.
-6 Discussion
-In this short paper, we contribute to our understanding of a crucial research tool.
-The behaviors we record offer both new strategies of working with the YouTube API, as well as opening further research directions to better understand how researchers can use it.
-6.1 Implications
-Our findings show that binning queries across the observation period is not as fruitful as previously thought [22, 24, 25], and offers low return on investment considering the quota cost of the search endpoint.
+6. Discussion
+In this short paper, we contribute to the understanding of a crucial research tool.
+The behaviors we record offer new strategies of working with the YouTube API and future research directions on expanding its use-cases.
+6.1. Implications
+Our findings show that binning queries across the observation period is not as fruitful as previously thought (Rieder et al., 2018; Violot et al., 2024; Yin and Brown, 2018), and offers low return on investment considering the quota cost of the search endpoint.
 Instead, researchers may experiment with breaking up their topics as opposed to their time frames.
@@
 The total number of results in the query metadata is a crucial way of assessing how optimal a query is (with lower being better/more stable).
-Moreover, researchers who use the API for data collection should aim to use ID-based endpoints wherever possible, as these are more complete and reliable.
-The strategy of pre-selecting channels, provided that there are ways to obtain an adequately representative sample, is a viable one as long as the search endpoint is not used to collect their videos.
-Instead, as we outline in Section 2, researchers can use a combination of the Channels: list and PlaylistItems: list endpoints, both of which take IDs as queries, to get complete channel uploads.
-Lastly, we urge researchers to clearly outline which endpoints they use in their data collection pipelines, as these choices can massively impact the replicability of their work.
-6.2 Limitations and Future Work
-The new strategies we put forward may need experimental validation, for example, by running progressively more restrictive queries and seeing how that influences the replicability of the data returned (alongside the reported video pool size).
-Moreover, we do not explicitly analyze how other parameters, such as ordering or including channel IDs, may affect replicability, although we have no reason to believe that this would alter our conclusions (especially as reverse chronological order should be immutable for historical queries, other than video deletions).
-Finally, there are several future directions that scholars can pursue to better understand this invaluable tool.
-For example, future research can replicate our experiments with more sparse collections over a longer period, to check for potential periodicity in set similarities.
-Moreover, given the substantial efforts that scholars have expended in creating sockpuppets for YouTube SERP audits [7, 8], future research can employ similar methods to ours to check the consistency between results of sockpuppet SERPs and search endpoint results.
+To reduce this number of results as much as possible, researchers should build their queries around API parameters that are specified in the documentation to affect search results (e.g., query, region code, etc.), and not merely the API response (e.g., before-after datetimes, topic IDs, etc.)
+Alternatively, in cases where discovery is possible at the channel level instead of using keywords, or where consistency takes precedence over data completeness, ID-based endpoints are a viable route.
+For example, as we outline in Section 2, a combination of the Channels: list and PlaylistItems: list endpoints, both of which take IDs as queries, would allow for the retrieval of complete channel uploads.
+We also urge researchers to specify these endpoints in their data collection pipelines, as these choices can massively impact the replicability of their work.
+6.2. Limitations and Future Work
+The strategy of using parameters that restrict the potential results pool is not explicitly quantified here, and can be addressed in future work that designs progressively restrictive queries using other parameters as we do here with time constraints.
+Moreover, we do not explicitly analyze how other parameters, such as ordering or including channel IDs, may affect replicability.
+Future work can also replicate our experiments with more sparse collections over a longer period, to check for potential periodicity in set similarities.
+Moreover, given the substantial efforts that scholars have expended in creating sockpuppets for YouTube SERP audits (Hussein et al., 2020; Jung et al., 2025), similar methods to ours can be employed to check the consistency between results of sockpuppet SERPs and search endpoint results.
 This would help us understand if the search endpoint has research value beyond data collection, for example, as a low-resource way of conducting SERP audits.
-Acknowledgements.
+Acknowledgments
 This work has been supported by the University of Washington’s Center for an Informed Public, the John S.
@@
 Knight Foundation (G-2019-58788), and the William and Flora Hewlett Foundation (2023-02789).
-The author is grateful to Adrian Padilla, Bernhard Rieder, and Ethan Zuckerman for feedback, comments, and discussions related to this work.
-Appendix A Query Parameters
+The author is grateful to Adrian Padilla, Bernhard Rieder, Ethan Zuckerman, and the anonymous reviewers for their helpful feedback.
+Appendix A Ethics
+This work makes sole use of video, comment, or channel IDs and high-level metadata (number of views, likes, etc.) as data points and does not analyze content beyond this point.
+The aim of this paper is to offer a better understanding of how data from YouTube, one of the largest platforms worldwide, should be understood and used by researchers.
+The societal benefits arising from this work apply insofar as the important topics studied on YouTube that we cover in the Introduction benefit from the directions we offer.
+Appendix B Query Parameters
 Our general query parameters, followed by topic-specific parameters (keywords and dates).
-A.1 General parameters
+B.1. General parameters
 Unless [variable], these parameters were kept consistent across queries.
@@
         "q": [variable]
-A.2 Topic-specific parameters
+B.2. Topic-specific parameters
 Keywords (q) and dates queried. Note that, for dates, we passed “publishedAfter” as the topic-specific date -14 and “publishedBefore” as +14 days.
@@
         "q": "fifa world cup"
-Appendix B ID-Based Queries
+Appendix C ID-Based Queries
 This section covers tests conducted with API endpoints that accept video or other IDs as queries. These endpoints show stable behavior and return mostly consistent data.
-B.1 Video: list Endpoint
+C.1. Video: list Endpoint
 In Figure 5, we show common video IDs between collections at a given time t and the previous time, as well as the first collection.
@@
 Moreover, the fact that we do not find consistent patterns between comparison ID and $J(S_{t},S_{1})$ suggests that API gaps in returning specific video metadata are not systematic, and are thus likely errors rather than intentional API behavior.
-Figure 4: Parallel plots of percentage of common videos retrieved between successive runs and overall Jaccard similarity of common videos retrieved using the “Videos: list” endpoint.
+Figure 5.
+Parallel plots of percentage of common videos retrieved between successive runs and overall Jaccard similarity of common videos retrieved using the “Videos: list” endpoint.
 Comparison ID refers to the respective collection (higher ID = later collection).
-B.2 CommentThreads: list and Comments: list Endpoints
+C.2. CommentThreads: list and Comments: list Endpoints
+topic TL, NS N, NS TL, S N, S
+BLM .329 .307 .976 .983
+Brexit .381 .339 .999 .999
+Capitol .648 .625 .998 .994
+Grammys .728 .737 .996 .992
+Higgs .974 N/A .998 N/A
+World Cup .470 .532 .999 .999
+Table 4.
+Jaccard similarities between first- and last-collection comment sets.
+TL = top-level, N = nested.
+NS = non-shared videos (full sets), S = shared videos only.
+N/A values for the Higgs topic, which is the oldest, are possibly due to the comment reply affordance operating differently in 2012.
 We also query the CommentThreads: list endpoint, which accepts video IDs as queries and returns all comment threads (with a maximum of five nested comments), as well as the Comments: list endpoint, which accepts thread IDs as queries and returns all nested comments.
@@
 With respect to comments drawn from common videos, differences between both nested and top-level comments between collections are negligible, showing that this endpoint itself does not systematically randomize results and is likely returning (almost) all comments for every queried video.
-topic TL, NS N, NS TL, S N, S
-BLM .329 .307 .976 .983
-Brexit .381 .339 .999 .999
-Capitol .648 .625 .998 .994
-Grammys .728 .737 .996 .992
-Higgs .974 N/A .998 N/A
-World Cup .470 .532 .999 .999
-Table 5: Jaccard similarities between first- and last-collection comment sets.
-TL = top-level, N = nested.
-NS = non-shared videos (full sets), S = shared videos only.
-N/A values for the Higgs topic, which is the oldest, are possibly due to the comment reply affordance operating differently in 2012.
-Appendix C Regression Robustness Checks
-Alternative regression model setups as robustness checks against the binned regression model implemented in the main paper.
-C.1 Frequency as Continuous Variable
-We use frequency as our dependent variable in a multiple Ordinary Least Squares (OLS) regression with robust standard errors.
-The overall model is significant ($F_{(14,5348)}=122.3,p<0.001$) and shows modest fit ($R^{2}=0.164$).
-We report standardized beta coefficients with confidence metrics in Table 6.
-The patterns are identical to what is reported in the main paper.
+Appendix D Regression Robustness Checks
+Alternative regression model setups as robustness checks against the SHAP-explained gradient boosting model implemented in the main paper.
+D.1. Binned Ordinal Regression
+We split the frequency of returns into four roughly equal bins (1-5, 6-10, 11-15, 16), taking into account that 16 is the modal value.
+We then perform an ordinal regression using a logit link function.
+This function is chosen due to the uniform distribution arising from the binning.
+All continuous features are log-transformed to reduce multicollinearity and standardized for better comparison between coefficients.
+A log-likelihood test shows that this model significantly outperforms the null model ($\chi^{2}=1137.63,p<0.001)$, although the overall fit is low (pseudo-$R^{2}=0.079$).
+This suggests either that several other factors may influence video appearance, or that much of the variance is indeed random.
+We show the standardized beta coefficients and confidence measures in Table 5.
+Topics are compared against BLM as the reference category, and the standard quality is compared against HD.
+Variable $\beta$ SE 95% CI
+SD (quality) -0.018 0.079 [-0.173, 0.137]
+brexit (topic) ***1.231 0.098 [1.039, 1.423]
+capriot (topic) -0.160 0.093 [-0.341, 0.022]
+grammys (topic) *0.171 0.083 [0.008, 0.333]
+higgs (topic) ***3.10 0.141 [2.826, 3.379]
+worldcup (topic) 0.161 0.101 [-0.037, 0.359]
+duration ***-0.115 0.028 [-0.170, -0.061]
+views 0.161 0.088 [-0.011, 0.333]
+likes **0.285 0.095 [0.098, 0.471]
+comments 0.069 0.064 [-0.058, 0.195]
+channel age 0.049 0.031 [-0.012, 0.110]
+channel views *0.3176 0.135 [0.053, 0.582]
+channel subs **-0.3784 0.122 [-0.617, -0.140]
+# channel videos -0.0212 0.075 [-0.169, 0.126]
+Table 5. Standardized regression coefficients for binned ordinal model. *$p<0.05$, **$p<0.01$, ***$p<0.001$.
+D.2. Frequency as Continuous Variable
 Variable $\beta$ SE 95% CI
@@
 # channel videos -0.2212 0.208 [-0.629, 0.187]
-Table 6: Standardized regression coefficients for OLS model. *$p<0.05$, **$p<0.01$, ***$p<0.001$.
-C.2 Non-Binned Ordinal Regression
+Table 6. Standardized regression coefficients for OLS model. p<0.05*p<0.05, p<0.01**p<0.01, p<0.001***p<0.001.
+We use frequency as our dependent variable in a multiple Ordinary Least Squares (OLS) regression with robust standard errors.
+The overall model is significant ($F_{(14,5348)}=122.3,p<0.001$) and shows modest fit ($R^{2}=0.164$).
+We report standardized beta coefficients with confidence metrics in Table 6.
+The patterns are identical to the previous binned ordinal regression model and to what is reported in the main paper.
+D.3. Non-Binned Ordinal Regression
 Ordinal regression where frequencies are treated as 16 distinct categories.
@@
 # channel videos -0.0958 0.049 [-0.193, 0.001]
-Table 7: Standardized regression coefficients for full ordered model. *$p<0.05$, **$p<0.01$, ***$p<0.001$.
+Table 7. Standardized regression coefficients for full ordered model. *$p<0.05$, **$p<0.01$, ***$p<0.001$.